# TAT-LLM: A Specialized Language Model for Discrete Reasoning over Tabular and Textual Data

### v1

@@ 4.1. Datasets
FinQA (Chen et al., 2021) 6,251 883 1,147
TAT-QA (Zhu et al., 2021) 13,215 1,668 1,669
TAT-DQA (Zhu et al., 2022) 13,251 1,645 1,662
4.2. Compared Models
We compare our TAT-LLM model with two kinds of models: fine-tuned models on the tabular and textual QA dataset, and LLMs in zero-shot setting. For fine-tuned models, we select the state-of-the-art supervised fine-tuned models for each dataset, which are listed separately in Table 2, Table 3 and Table 4. For LLMs, we utilize the state-of-the-art GPT3.5-Turbo (Brown et al., 2020) and GPT-4 (OpenAI, 2023). Besides, we also adopt powerful smaller LLMs that have garnered considerable research interest in the field, including Vicuna (7B) (Chiang et al., 2023) and LLaMA 2-Chat (7B and 70B) (Touvron et al., 2023b). In addition to these general LLMs, we also compare our models with LLMs specialized in math word problems, including MAmmoTH (70B) (Yue et al., 2023) and WizardMath (70B) (Luo et al., 2023), since our task involves numerical reasoning. All LLMs are tested in zero-shot setting, because financial tabular and textual data inputs tend to be lengthy, making it impractical to include additional in-context examples due to input length limits.
4.3. Implementation Details
Since the question type and the scale of the answer have been annotated in both TAT-QA and TAT-DQA datasets, we add two more steps in the instruction part of the input instance, in addition to the three steps in the proposed Step-wise Pipeline shown in Figure 2 c). One step is Question Type Predictor, with which we limit the model to select one value from the following question types: Span, Multiple Spans, Counting, and Arithmetic. For Span and Multiple Spans questions, we set the output of the Reasoner to “N.A.”. The other step is Scale Predictor, with which we restrict the model to choose one from the following values: Thousand, Million, Billion, Percent and None. For all the templates used to construct training instances of our  TAT-LLM model, please refer to Appendix A.
We train our TAT-LLM model by fine-tuning LLaMA 2 (Touvron et al., 2023b) using Low-Rank Adaptation (LoRA) (Hu et al., 2022) on one NVIDIA DGX-A100 with eight A100 GPUs. The quantization is 8 bit. We use Adam optimizer with a learning rate of $43e-$ 4 and warmup over the first $3\%$ steps to train. The maximum sequence length is $4,096$ and the maximum number of epochs is set to $3$. The batch size is set to $4$ and the gradient accumulation step is $10$.
For comparing with fine-tuned models, we take the results from their original papers respectively. For each LLM, we utilize three different templates to perform zero-shot inference to obtain prediction results. We select the best result as the reported result in Section 4.5. Please refer to Appendix B for details of the templates we use for zero-shot inference. We utilize the latest version [^1] Date: Sep 2023 GPT3.5-Turbo (Brown et al., 2020) and GPT-4 (OpenAI, 2023) via OpenAI APIs [^2] https://platform.openai.com/.. We set the temperature as 0, top p as 1.0, max token as $1,000$, and other parameters as default. We obtain the official trained checkpoint of Vicuna (Chiang et al., 2023), LLaMA 2-Chat (Touvron et al., 2023b), MAmmoTH (Yue et al., 2023) and WizardMath (Luo et al., 2023) from Hugginface [^3] https://huggingface.co/models. The inference is done on one NVIDIA DGX-A100 with eight A100 GPUs. The parameters num_beam and do_sample are 1 and false respectively.
4.4. Evaluation Metrics
We adopt the standard automatic evaluation metrics of FinQA, TAT-QA and TAT-DQA to evaluate the compared methods. For TAT-QA and TAT-DQA, we adopt the Exact Match (EM) and the numeracy-focused (macro-averaged) F1 score (Zhu et al., 2021, 2022). Both two metrics measure the overlap between a bag-of-words representation of the gold and predicted answers. The numeracy-focused F1 score is set to 0 unless the predicted number is exactly equal to the ground truth. We use EM for FinQA, which is the same as the metric of Execution Accuracy originally used in FinQA (Chen et al., 2021).
For zero-shot prediction using LLMs like GPT3.5-Turbo (Brown et al., 2020) and WizardMath (Luo et al., 2023), we omit the scale prediction and compare the predicted value with the ground truth value only, assuming the scale prediction is always correct. For our TAT-LLM model, since we instruct the model to output a structured table that is friendly to automatic evaluation, we take the scale into account.
4.5. Main Results
Table 2. Performance of our TAT-LLM model and compared models on the test set of FinQA. The best result is marked in bold and the second-best result is underlined. The number in red indicates the improvement over the second-best result.
@@ 4.5. Main Results
Fine-tuned Longformer (Beltagy et al., 2020) 21.90
NeRd (Chen et al., 2019) 48.57
[…]
Zero-shot Vicuna (7B) (Chiang et al., 2023) 10.11
LLaMA 2-Chat (7B)  (Touvron et al., 2023b) 15.43
LLaMA 2-Chat (70B)  (Touvron et al., 2023b) 32.17
MAmmoTH (70B) (Yue et al., 2023) 36.09
WizardMath (70B) (Luo et al., 2023) 47.25
GPT-3.5-Turbo (Brown et al., 2020) 58.00
GPT-4 (OpenAI, 2023) 63.91
Ours TAT-LLM (7B) 64.60 (+0.69)
We first compare the performance of our TAT-LLM  with compared methods on each tabular and textual QA dataset, i.e., FinQA, TAT-QA and TAT-DQA, respectively. The experimental results are summarized in Table 2 for FinQA, Table 3 for TAT-QA and Table 4 for TAT-DQA. From the tables, we make the following observations. 1) Our TAT-LLM (7B) significantly outperforms all other models on all three datasets, including the previous best fine-tuned models and the state-of-the-art GPT-4 (OpenAI, 2023). In particular, our TAT-LLM reaches $64.60\%$, $74.56\%$ and $69.45\%$ in terms of EM on the test set of FinQA, TAT-QA and TAT-DQA respectively, i.e. an increase of $0.69$, $2.64$ and $4.99$ points compared to GPT-4. These results well demonstrate the great effectiveness of our model. The results also confirm the rousing potential of specializing smaller language models, like our TAT-LLM (7B) model, for specific tasks to yield better performance than very large-scale models like GPT-4. 2) GPT-4 with our Step-wise Pipeline surpasses the best fine-tuned models on all datasets, showing the rationality of such a pipeline. 3) However, the performance of GPT-4 and our TAT-LLM  obviously lags behind that of human experts, showing that this task is still challenging. 4) The strong general LLM GPT3.5-Turbo and LLaMA 2-Chat (70B) underperform the best fine-tuned models, evidencing that supervised fine-tuning is still essential for achieving advanced performance on table-text QA datasets. 5) The LLMs specialized in math word problems, i.e., WizardMath (Luo et al., 2023) and MAmmoTH (Yue et al., 2023), largely underperform GPT3.5-Turbo, GPT-4 and our TAT-LLM model, showing that current numerically-enhanced LLMs still struggle in discrete reasoning over tabular and textual QA. This again speaks for the considerable value of specializing models for different tasks.
Table 3. Performance of our TAT-LLM model and compared models on the test set of TAT-QA. Best results are marked in bold and second-best results are underlined. Numbers in red indicate the improvement over the second-best results.
@@
Fine-tuned TagOp (Zhu et al., 2021) 50.10 58.00
TeaBReaC (Trivedi et al., 2022) 55.80 63.80
KIQA (Nararatwong et al., 2022b) 58.20 67.40
FinMath (Li et al., 2022b) 58.30 68.20
GANO (Nararatwong et al., 2022a) 61.90 72.10
MHST (Zhu et al., 2022) 63.60 72.70
UniPCQA (Deng et al., 2022) 63.90 72.20
SoarGraph (Zhu et al., 2023b) 65.40 75.30
UniRPG (Zhou et al., 2022) 67.20 76.00
RegHNT (Lei et al., 2022) 70.30 77.90
MVGE (Wei et al., 2023) 70.90 79.10
Zero-shot Vicuna (7B) (Chiang et al., 2023) 32.53 40.97
LLaMA 2-Chat (7B) (Touvron et al., 2023b) 37.16 45.37
MAmmoTH (70B) (Yue et al., 2023) 38.97 46.51
WizardMath (70B) (Luo et al., 2023) 39.63 45.28
LLaMA 2-Chat (70B) (Touvron et al., 2023b) 45.94 53.80
GPT-3.5-Turbo (Brown et al., 2020) 59.47 68.11
GPT-4 (OpenAI, 2023) 71.92 79.71
Ours TAT-LLM (7B) 74.56 (+2.64) 82.88 (+3.17)
4.6. Model Analysis
4.6.1. Fine-tuning with Combined Data.
[…]
Table 4. Performance of our TAT-LLM  model and compared models on the test set of TAT-DQA. Best results are marked in bold and second-best results are underlined. Numbers in red indicate the improvement over the second-best results.
@@ 4.6. Model Analysis
Fine-tuned NumNet+ V2 (Ran et al., 2019) 30.60 40.10
TagOp (Zhu et al., 2021) 33.70 42.50
MHST (Zhu et al., 2022) 41.50 50.70
Doc2SoarGraph (Zhu et al., 2023b) 59.20 67.60
Zero-shot Vicuna (7B) (Chiang et al., 2023) 28.44 36.72
LLaMA 2-Chat (7B)  (Touvron et al., 2023b) 34.52 42.32
MAmmoTH (70B) (Yue et al., 2023) 35.42 42.82
WizardMath (70B) (Luo et al., 2023) 36.44 41.55
LLaMA 2-Chat (70B)  (Touvron et al., 2023b) 41.91 49.74
GPT-3.5-Turbo (Brown et al., 2020) 52.74 61.40
GPT-4 (OpenAI, 2023) 64.46 72.20
Ours TAT-LLM (7B) 69.45 (+4.99) 77.75 (+5.55)
Table 5. Performance of the TAT-LLM model on the test sets, which is trained using a combination of all three training sets. Best results are marked in bold. Numbers in red indicate the improvement compared with GPT-4.
@@
4.6.2. Different Fine-tuning Strategies.
@@
  gains (+6.71) (+4.81) (+4.94) (+4.81) (+4.91)
GPT-4 63.91 71.92 79.71 64.46 72.20
To verify the effect of our fine-tuning strategy, i.e. Step-wise Pipeline plus External Executor, we compare it with separately applying End-to-end Pipeline or Step-wise Pipeline on LLaMA 2 (7B) models. The results are summarized in Figure 3. Firstly, we can see our TAT-LLM model clearly outperforms the two variants following the End-to-end Pipeline and the Step-wise Pipeline on all tabular and textual QA datasets for both evaluation metrics, demonstrating the effectiveness of our fine-tuning strategy. Besides, we find that the Step-wise Pipeline also outperforms the End-to-end Pipeline in all situations, showing the rationality of incorporating intermediate reasoning steps, which is in line with previous research on chain-of-thoughts reasoning (Wei et al., 2022c; Chung et al., 2022). Moreover, it is observed that the End-to-end Pipeline achieves notably poor performance on FinQA compared to that on TAT-QA and TAT-DQA. This is probably because FinQA requires more complex discrete reasoning steps than TAT-QA and TAT-DQA, and omitting such steps in fining-tuning would largely degrade model performance. Finally, we find a significant performance increase from the Step-wise Pipeline to TAT-LLM. This is due to the incorporation of the reliable External Executor which largely guarantees the reasoning accuracy of the answer. In the next section, we perform detailed analysis on the effectiveness of the External Executor.
Figure 3. Comparison on different fine-tuning strategies.
An example of QA with numerical reasoning over financial tabular and textual data
Figure 3. Comparison on different fine-tuning strategies.
4.6.3. Effect of External Executor w.r.t. Different Model Sizes.
@@
Figure 4. Performance comparison in terms of EM between TAT-LLM (7B) and LLaMA 2-Chat (7B) for different question types on TAT-QA. Numbers in red denote the relative improvement from LLaMA 2-Chat (7B) to TAT-LLM (7B).
An example of QA with numerical reasoning over financial tabular and textual data
@@
Figure 4. Performance comparison in terms of EM between TAT-LLM (7B) and LLaMA 2-Chat (7B) for different question types on TAT-QA. Numbers in red denote the relative improvement from LLaMA 2-Chat (7B) to TAT-LLM (7B).
Figure 5. Performance comparison in terms of EM between TAT-LLM (7B) and LLaMA 2-Chat (7B) for different question types on TAT-DQA. Numbers in red denote the relative improvement from LLaMA 2-Chat (7B) to TAT-LLM (7B).
4.6.4. Improvements w.r.t. Different Question Types.
@@
Here we analyze the improvements achieved by TAT-LLM model regarding different question types of TAT-QA and TAT-DQA datasets. We compare the performance of TAT-LLM (7B) and LLaMA 2-Chat (7B) for each question type within TAT-QA and TAT-DQA. The results are summarized in Figure 4 and Figure 5 for TAT-QA and TAT-DQA, respectively. We can observe substantial performance improvements for the Arithmetic questions, which are over two-fold, for both datasets. The larger performance gains are attributed to the effective fine-tuning strategy of TAT-LLM  as well as the reliable External Executor that ensures accurate execution of the equations and logic rules. Large performance improvements are also found for the Counting questions, namely 109.1% for TAT-QA and 53.8% for TAT-DQA. This is probably due to executing more correct counting operations. For the Span and Multiple Spans types, relatively smaller improvements are observed, which are still over 30%. This is potentially because these questions involve simpler or no discrete reasoning, mainly numerical comparison.
Table 7. Examples of errors on the test set of FinQA and corresponding percentages in each step. Q, G and P denote question, ground truth and prediction.

### v3

@@ 3.2 TAT-LLM
FinQA 6,251 883 1,147
TAT-QA 13,215 1,668 1,669
TAT-DQA 13,251 1,645 1,662
Table 1: Statistics of the train, validation and test splits for the financial tabular and textual QA datasets.
[…]
To train the model, the response part in the training set needs to be generated in advance. Existing tabular and textual QA datasets often provide the answer to the question and the derivation annotated by human experts. To generate the correct response for constructing our training instance, we first automatically identify the supporting evidence from the derivation of each question, such as the numbers and entities, which are used as the output of the first step. We then convert the derivation to a valid equation or logic rule as the output of the second step, making sure its execution result is consistent with the final answer. We directly use the annotated answer as the final answer, which is also the output of the last step.
Training.
After constructing the training data, we train our TAT-LLM model in various sizes, including 7B, 13B, and 70B, by fine-tuning LLaMA 2 Touvron et al. (2023b) using Low-Rank Adaptation (LoRA) Hu et al. (2022). More details are provided in Appendix A.
External Executor.
We observe that the trained model struggles in performing the execution of the mathematical equations and logic rules (e.g., counting, comparison) in the last step, according to experimental evaluations in Section 4.3. Such incompetence of the Executor has also been found in previous works Luo et al. (2023); Yue et al. (2023). To enhance the model’s accuracy of the final output, we propose to add an External Executor to our model, which takes the model’s intermediate outputs as input and performs the execution of the equations or logic rules to obtain the final answer. With this External Executor, the model can refine its output to a better one. With the output of the model formatted to a structured table, we can easily access the intermediate results. After obtaining the intermediate results, the External Executor is applied to refine the final answer instead of directly using the prediction of the model. The whole process is summarized in Algorithm 1.
4 Experiments
@@
Fine-tuned Longformer 21.90
NeRd 48.57
$\text{FinQANet}_{\textit{BERT}}$ 50.00
$\text{DyRRen}_{\textit{BERT}}$ 59.37
$\text{FinQANet}_{\textit{RoBERTa}}$ 61.24
$\text{ELASTIC}_{\textit{RoBERTa}}$ 62.66
$\text{DyRRen}_{\textit{RoBERTa}}$ 63.30
Zero-shot Vicuna (7B) 10.11
LLaMA 2-Chat (7B) 15.43
LLaMA 2-Chat (70B) 32.17
MAmmoTH (70B) 36.09
WizardMath (70B) 47.25
GPT3.5-Turbo 58.00
GPT-4 63.91
Ours TAT-LLM (7B) ( +0.69)
64.60
Table 2: Performance of our TAT-LLM model and compared models on the test set of FinQA. Best results are marked in bold and numbers in red indicate the improvement over the underlined second-best results.
4.1 Datasets, Models and Evaluation Metrics
Datasets.
We use FinQA  Chen et al. (2021), TAT-QA Zhu et al. (2021) and TAT-DQA Zhu et al. (2022) for our experiments. See Table 1 for the statistics of splits of each dataset.
$\bullet$ FinQA Chen et al. (2021) is an expert-annotated tabular and textual QA dataset in which the tables and text are sampled from financial reports. It focuses on discrete reasoning capabilities like addition, subtraction, multiplication, division, and numerical comparison.
$\bullet$ TAT-QA Zhu et al. (2021) is also built with tables and paragraphs extracted from financial reports. Most questions require discrete reasoning to generate the answers, and meanwhile, there are also cases where the answers can be extracted directly from the tables or text. Its questions are classified into four different types: Span, Multiple Spans, Counting, and Arithmetic.
$\bullet$ TAT-DQA Zhu et al. (2022) is an extension of the TAT-QA dataset, focusing on question answering over the original long financial statements with up to three pages, and the position and structure of the tables are unknown.
@@ 4.1 Datasets, Models and Evaluation Metrics
Fine-tuned TagOp 50.10 58.00
TeaBReaC 55.80 63.80
KIQA 58.20 67.40
FinMath 58.30 68.20
GANO 61.90 72.10
MHST 63.60 72.70
UniPCQA 63.90 72.20
SoarGraph 65.40 75.30
UniRPG 67.20 76.00
RegHNT 70.30 77.90
MVGE 70.90 79.10
Zero-shot Vicuna (7B) 32.53 40.97
LLaMA 2-Chat (7B) 37.16 45.37
MAmmoTH (70B) 38.97 46.51
WizardMath (70B) 39.63 45.28
LLaMA 2-Chat (70B) 45.94 53.80
GPT3.5-Turbo 59.47 68.11
GPT-4 71.92 79.71
Ours TAT-LLM (7B) (+2.64) (+3.17)
74.56 82.88
Table 3: Performance of our TAT-LLM model and compared models on the test set of TAT-QA.
Compared Models.
We compare our TAT-LLM model with two kinds of models: fine-tuned models on the tabular and textual QA dataset, and LLMs in zero-shot setting. For fine-tuned models, we select the state-of-the-art supervised fine-tuned models for each dataset, which are listed separately in Table 2, Table 3 and Table 4. For LLMs, we utilize the state-of-the-art GPT3.5-Turbo Brown et al. (2020) and GPT-4 OpenAI (2023). Besides, we also adopt powerful smaller LLMs that have garnered considerable research interest in the field, including Vicuna (7B) Chiang et al. (2023) and LLaMA 2-Chat (7B and 70B) Touvron et al. (2023b). In addition to these general LLMs, we also compare our models with LLMs specialized in math word problems, including MAmmoTH (70B) Yue et al. (2023) and WizardMath (70B) Luo et al. (2023), since our task involves numerical reasoning. All LLMs are tested in zero-shot setting, because financial tabular and textual data inputs tend to be lengthy, making it impractical to include additional in-context examples due to input length limits.
Evaluation Metrics.
For TAT-QA and TAT-DQA datasets, we adopt the Exact Match (EM) and the numeracy-focused (macro-averaged) F1 score Zhu et al. (2021, 2022). Both two metrics measure the overlap between a bag-of-words representation of the gold and predicted answers. The numeracy-focused F1 score is set to 0 unless the predicted number is exactly equal to the ground truth. We use EM for FinQA, which is the same as the metric of Execution Accuracy originally used in FinQA Chen et al. (2021).
For zero-shot prediction using LLMs, we omit the scale prediction and compare the predicted value with the ground truth value only, assuming the scale prediction is always correct. For our TAT-LLM model, since we instruct the model to output a structured table that is friendly to automatic evaluation, we take the scale into account.
@@
Fine-tuned NumNet+ V2 30.60 40.10
TagOp 33.70 42.50
MHST 41.50 50.70
Doc2SoarGraph 59.20 67.60
Zero-shot Vicuna (7B) 28.44 36.72
LLaMA 2-Chat (7B) 34.52 42.32
MAmmoTH (70B) 35.42 42.82
WizardMath (70B) 36.44 41.55
LLaMA 2-Chat (70B) 41.91 49.74
GPT3.5-Turbo 52.74 61.40
GPT-4 64.46 72.20
Ours TAT-LLM (7B) ( +4.99) (+5.55)
69.45 77.75
Table 4: Performance of our TAT-LLM  model and compared models on the test set of TAT-DQA.
4.2 Main Results
We first compare the performance of our TAT-LLM  with previous methods on each dataset respectively. The experimental results are summarized in Table 2 for FinQA, Table 3 for TAT-QA and Table 4 for TAT-DQA. From the tables, we make the following observations. 1) Our TAT-LLM (7B) significantly outperforms all the previous models on each of the three datasets, including the previous best fine-tuned models and the state-of-the-art GPT-4 OpenAI (2023). In particular, our TAT-LLM reaches $64.60\%$, $74.56\%$ and $69.45\%$ in terms of EM on the test set of FinQA, TAT-QA and TAT-DQA respectively, i.e. an increase of $0.69$, $2.64$ and $4.99$ points compared to GPT-4. These results well demonstrate the noticeable effectiveness of our TAT-LLM model. The results also confirm the rousing potential of specializing smaller language models, like our TAT-LLM (7B) model, for specific tasks to yield better performance than very large-scale models like GPT-4. 2) However, the performance of GPT-4 and our TAT-LLM (7B) obviously lags behind that of human experts, showing that this task is still challenging. 3) The strong general LLM GPT3.5-Turbo Ouyang et al. (2022) and LLaMA 2-Chat (70B) Touvron et al. (2023b) underperform the best fine-tuned models, evidencing that supervised fine-tuning is still essential for achieving advanced performance for this challenge. 4) The LLMs specialized in mathematical reasoning, i.e., WizardMath Luo et al. (2023) and MAmmoTH Yue et al. (2023), largely underperform GPT3.5-Turbo, GPT-4 and our TAT-LLM model, showing that current numerically-enhanced LLMs still struggle in discrete reasoning over tabular and textual QA. This again speaks for the considerable value of specializing models for specific tasks.
@@ 4.2 Main Results
Table 5: Performance of the TAT-LLM model trained with combination of all three training sets.
4.3 In-Depth Analysis
Fine-tuning with Combined Data.
@@ 4.3 In-Depth Analysis
gains (+) 6.71 4.81 4.94 4.81 4.91
@@
Figure 3: Comparison of different training strategies.
Different Fine-tuning Strategies.
To verify the effect of our fine-tuning strategy, i.e. Step-wise Pipeline plus External Executor, we compare it with separately applying End-to-end Pipeline or Step-wise Pipeline on LLaMA 2 (7B) models. The results are summarized in Figure 3. Firstly, we can see our TAT-LLM (7B) clearly outperforms the two variants following the End-to-end Pipeline and the Step-wise Pipeline on all tabular and textual QA datasets for both evaluation metrics, demonstrating the effectiveness of our fine-tuning strategy. Besides, we find that the Step-wise Pipeline also outperforms the End-to-end Pipeline in all situations, showing the rationality of incorporating intermediate reasoning steps, which is in line with previous research on chain-of-thoughts reasoning Wei et al. (2022c). Moreover, it is observed that the End-to-end Pipeline achieves notably poor performance on FinQA compared to that on TAT-QA and TAT-DQA. This is probably because FinQA requires more complex discrete reasoning steps than TAT-QA and TAT-DQA, and omitting such steps in fining-tuning would largely degrade model performance. Finally, we find a significant performance increase from the Step-wise Pipeline to TAT-LLM. This is due to the incorporation of the reliable External Executor which largely guarantees the reasoning accuracy of the answer. In the next section, we perform detailed analysis on the effectiveness of the External Executor.
@@
Figure 4: Performance comparison in terms of EM between TAT-LLM (7B) and LLaMA 2-Chat (7B) for different question types on TAT-QA.
Effect of External Executor w.r.t. Different Model Sizes.
[…]
@@
Figure 5: Performance comparison in terms of EM between TAT-LLM (7B) and LLaMA 2-Chat (7B) for different question types on TAT-DQA.
@@
Improvements w.r.t. Different Question Types. Here we analyze the improvements achieved by TAT-LLM model regarding different question types of TAT-QA and TAT-DQA datasets. We compare the performance of TAT-LLM (7B) and LLaMA 2-Chat (7B) for each question type. The results are summarized in Figure 4 and Figure 5 for TAT-QA and TAT-DQA, respectively. We can observe substantial performance improvements for the Arithmetic questions, which are over two-fold, for both datasets. The larger performance gains are attributed to the effective fine-tuning strategy of TAT-LLM  as well as the reliable External Executor that ensures accurate execution of the equations and logic rules. Large performance improvements are also found for the Counting questions, namely 109.1% for TAT-QA and 53.8% for TAT-DQA. This is probably due to executing more correct counting operations. For the Span and Multiple Spans types, relatively smaller improvements are observed, which are still over 30%. This is potentially because these questions involve simpler or no discrete reasoning, mainly numerical comparison.
5 Conclusion
In this work, we first abstract a Step-wise Pipeline for tabular and textual QA based on the previous multi-step methods. Following this pipeline, we develop TAT-LLM model by specializing smaller language models (i.e., 7B) in discrete reasoning over tabular and textual data. We validate its effectiveness with extensive experiments, showing that our TAT-LLM model can outperform both conventional methods and very large-scale LMs like GPT-4 on this task. Our work well demonstrates that specialized models might be a promising direction towards more advanced models on specific tasks that can compete with human experts.
Limitations
Despite the impressive performance on all three datasets i.e., FinQA Chen et al. (2021), TAT-QA Zhu et al. (2021), and TAT-DQA Zhu et al. (2022), our TAT-LLM model still has much room for further improvement, as shown in error analysis in Appendix B. In addition, our TAT-LLM model is designed for the documents that contain both tabular and textual data and feature rich numerical values. This means it may have limited advantages over other kinds of documents like pure textual documents. Also, our model may not be directly applied to understand the documents with a large number of pages (e.g., >100 pages) due to the constraint of the maximum input sequence length.
Ethics Statement
In this work, we first abstract a Step-wise Pipeline for tabular and textual QA. Then, we propose the TAT-LLM model by specializing a smaller language model (i.e., LLaMA 2 7B, 13B and 70B) in discrete reasoning over a hybrid of tabular and textual data. Our TAT-LLM model is developed on open-source tools and datasets to assist human beings in processing and understanding such kind of data. Thus, we do not anticipate any potential risks or negative ethical issues.
Appendix A Implementation Details
Since the question type and the scale of the answer have been annotated in both TAT-QA and TAT-DQA datasets, we add two more steps in the instruction part of the input instance, in addition to the three steps in the proposed Step-wise Pipeline shown in Figure 2 c). One step is Question Type Predictor, with which we limit the model to select one value from the following question types: Span, Multiple Spans, Counting, and Arithmetic. For Span and Multiple Spans questions, we set the output of the Reasoner to “N.A.”. The other step is Scale Predictor, with which we restrict the model to choose one from the following values: Thousand, Million, Billion, Percent and None. For all the templates used to construct training instances of our  TAT-LLM model, please refer to Appendix C.
We train our TAT-LLM model on one NVIDIA DGX-A100 with eight A100 GPUs. The quantization is 8 bit. We use Adam optimizer with a learning rate of $43e-$ 4 and warmup over the first $3\%$ steps to train. The maximum sequence length is $4,096$ and the maximum number of epochs is set to $3$. The batch size is set to $4$ and the gradient accumulation step is $10$.
For comparing with fine-tuned models, we take the results from their original papers respectively. For each LLM, we utilize three different templates to perform zero-shot inference to obtain prediction results. We select the best result as the reported result in Section 4.2. Please refer to Appendix D for details of the templates we use for zero-shot inference. We utilize the latest version [^1] Date: Sep 2023 GPT3.5-Turbo Ouyang et al. (2022) and GPT-4 OpenAI (2023) via OpenAI APIs. We set the temperature as 0, top p as 1.0, max token as $1,000$, and other parameters as default. We obtain the official trained checkpoint of Vicuna Chiang et al. (2023), LLaMA 2-Chat Touvron et al. (2023b), MAmmoTH Yue et al. (2023) and WizardMath Luo et al. (2023) from Hugginface. The inference is done on one NVIDIA DGX-A100 with eight A100 GPUs. The parameters num_beam and do_sample are 1 and false respectively.